OpenStudy (anonymous):

(7 × 10^6) × (6 × 10^-4) How can I simplify this? I don't want the answer, I would appreciate the steps for future reference :]

OpenStudy (bibby):

you can rearrange the numbers using the properties of multiplication. for example \((a \times b)*(c \times d)=(a \times c)*(b \times d)\) if you want steps, youtube or google "multiplying scientific notation"

OpenStudy (bibby):

also recall when multiplying exponents with a common base, you just add the powers

OpenStudy (anonymous):

is it 13x10^2?

OpenStudy (bibby):

the 6 and 7 are still being multiplied like they normally would be

OpenStudy (anonymous):

i added the bases ( 7+6) to get 13

OpenStudy (bibby):

when I said add the powers I meant the powers. For example, \(3^6*3^4=3^{6+4}=3^{10}\)

OpenStudy (bibby):

so \((7*6)(10^6*10^{-4})\)

OpenStudy (anonymous):

I have options: 4.2x10^3 4.2x10^1 13x10^2 4.1x10^4

OpenStudy (anonymous):

but i know but a fact that 7 times 6 is 42 so 4.2?

OpenStudy (bibby):

you already simplified the rightmost parentheses \(\large (7 * 6)*10^2\)

OpenStudy (bibby):

yeah but we have to adjust the exponent

OpenStudy (bibby):

\(\huge 42*10^2=4.2*10^?\)

OpenStudy (anonymous):

2?

OpenStudy (anonymous):

im sort of lost sorry

OpenStudy (bibby):

yeah, this is the part I kinda suck at explaining, which is why I suggested youtube As I understand it, we're moving the decimal point once to the left, so we need to account for it by multiplying the other side with a 10 and \(10^2*10=10^{(2+1)}=10^3\)

OpenStudy (bibby):

if you want, in this case you can just simplify the numbers.  rofl it's so hard doing this on a laptop